# An Intention-driven Lane Change Framework Considering Heterogeneous Dynamic Cooperation in Mixed-traffic Environment

### v2

@@ 5.2 Main Results
5.3 Ablation Study
Table 3 presents the ablation results of the proposed framework by selectively removing different core and cooperation modules. Using only the BC module achieves an F1-score of 0.9240, which indicates that pure behavior cloning provides a reasonable baseline but lacks robustness. Adding IRL alone degrades performance (F1-score 0.8939), suggesting that reward inference without supervised guidance may be unstable. Incorporating either the intrinsic cooperation score (LCS) or the dynamic cooperation score (DCS) individually improves the results (F1-scores 0.9296 and 0.9249, respectively), highlighting their complementary roles in modeling driver heterogeneity. The full model, which integrates BC, IRL, LCS, and DCS, achieves the best performance across all metrics, with an accuracy of 0.9367 and an F1-score of 0.9379. These findings demonstrate that BC and IRL are synergistic, where BC stabilizes learning while IRL enhances reward-driven generalization, and cooperation-aware modules (LCS and DCS) provide additional discriminative power by explicitly capturing intrinsic driver styles and interactive dynamics.
5.4 Convergence Analysis of Cooperation-Aware Intention Module
Table 3: Ablation study results on the proposed framework.
@@ 5.5 Case Study
Figure 7 shows prediction snapshots at one-second intervals (10 steps) for the T-Rear vehicle. The left column illustrates our IRL-based method, while the right column shows IDM predictions. The results demonstrate that our method enables the ego vehicle to complete the lane change substantially faster and more efficiently. Specifically, in Fig. 7(a), the maneuver is completed within 4.6 seconds, whereas in Fig. 7(b), IDM requires 7.8 seconds and only reaches mid-transition at 4.6 seconds.
@@
6 Discussions and Conclusion
6.1 Limitations and Future Work
@@ 6.2 Conclusion
Acknowledgments
This work was partially supported by NSFC Grant U24A20252 and 62373315, Guangdong provincial project 2023ZT10X009, Nansha Key Science and Technology Project under Grant 2023ZD006, and Guangzhou Bureau Of Education College Scientific Research Project 2024312169.

### v3

@@ V-A Hyperparameter and Data Settings
V-B Main Results
  • IDM+MOBIL: A rule-based baseline that integrates the Intelligent Driver Model (IDM) for longitudinal control [42] with the MOBIL framework for lateral maneuvers [13]. Lane change decisions are triggered by predefined safety constraints and acceleration incentives, aiming to balance efficiency and safety.
  • H-LSTM: A learning-based approach employing a hierarchical long short-term memory (H-LSTM) network [43] to capture temporal dependencies in multi-vehicle interactions and enable online lane change prediction.
  • BC: A supervised imitation-learning baseline [44], where the policy directly maps state features to expert actions via behavior cloning, without explicit reward modeling.
  • VWC: A learning-based baseline employing a variable-weight cooperation (VWC) model [45], which adaptively assigns dynamic weights to lane change conditions and driving demands across different phases of interaction.
Evaluation metrics for lane change intention prediction include precision, recall, and the F1 score, defined as follows [46]:
precision $=\frac{TP}{TP+FP}$ (20)
recall $=\frac{TP}{TP+FN}$ (21)
Table II compares the performance of different baselines against our proposed method on LK and LC classification. The results clearly show that rule-based IDM+MOBIL performs the worst, with overall accuracy of only 0.6185, reflecting its limited adaptability to heterogeneous human driving patterns. Learning-based methods achieve significantly higher performance: H-LSTM reaches above 0.91 across all metrics by capturing temporal dependencies, while BC attains moderate performance (accuracy 0.8461) but is constrained by its reliance on direct imitation. The VWC model further improves classification accuracy to 0.9207 by adaptively weighting interactive features. In contrast, our proposed intention-driven BC-IRL framework achieves the best overall results, with accuracy 0.9418 and F1-score 0.9427. Notably, it improves lane change recognition (precision 0.8911, recall 0.9239, F1-score 0.9072) over all baselines, highlighting its ability to model inter-driver heterogeneity through cooperation scores and reward inference.
TABLE III: Ablation study results on the proposed framework.
@@ V-B Main Results
V-C Ablation Study
Table III presents the ablation results of the proposed framework by selectively removing different core and cooperation modules. Using only the BC module achieves an F1-score of 0.9240, which indicates that pure behavior cloning provides a reasonable baseline but lacks robustness. Adding IRL alone degrades performance (F1-score 0.8939), suggesting that reward inference without supervised guidance may be unstable. Incorporating either the intrinsic cooperation score (LCS) or the dynamic cooperation score (DCS) individually improves the results (F1-scores 0.9296 and 0.9249, respectively), highlighting their complementary roles in modeling driver heterogeneity. The full model, which integrates BC, IRL, LCS, and DCS, achieves the best performance across all metrics, with an accuracy of 0.9367 and an F1-score of 0.9379. These findings demonstrate that BC and IRL are synergistic, where BC stabilizes learning while IRL enhances reward-driven generalization, and cooperation-aware modules (LCS and DCS) provide additional discriminative power by explicitly capturing intrinsic driver styles and interactive dynamics.
V-D Convergence Analysis of Cooperation-Aware Intention Module
@@ V-E Case Study
Fig. 7 shows prediction snapshots at one-second intervals (10 steps) for the T-Rear vehicle. The left column illustrates our IRL-based method, while the right column shows IDM predictions. The results demonstrate that our method enables the ego vehicle to complete the lane change substantially faster and more efficiently. Specifically, in Fig. 7(a), the maneuver is completed within 4.6 seconds, whereas in Fig. 7(b), IDM requires 7.8 seconds and only reaches mid-transition at 4.6 seconds.
@@
VI Discussions and Conclusion
VI-A Limitations and Future Work